# Forbidding just one intersection for short integer sequences

### v1

@@ Abstract
In this paper, we study the famous Erdős–Sós forbidden intersection problem for words over alphabet of size $m$: what is the maximal size of a subfamily $\mathcal{F}$ of $m]^{n}$ that does not contain two vectors $x,y$ that coincide on exactly t1t-1 coordinates? We answer this question provided $m\geqslant\operatorname{poly}(t)$, $n\geqslant\operatorname{poly}(t)$ for some polynomial function $\operatorname{poly}(\cdot)$ of $t$, greatly extending the recent result of Keevash, Lifshitz, Long and Minzer. Our proof combines some of the recently developed methods in extremal combinatorics, including the spread approximation technique of Kupavskii and Zakharov, and the hypercontractivity approach developed in a series of works of Keevash, Keller, Lifshitz, Long, Marcus and Minzer.
@@ 1 Introduction
Let $\binom{[n]}{k}$ be a family of $k$-subsets of $n]=\{1,\ldots,n\}$. In 1971, Erdős and Sós (see [erdos1975problems]) considered the following problem.
@@
Problem 1 is a harder version of the following question raised by Erdős, Ko and Rado [erdos1961intersection] ten years earlier.
@@
Erdős, Ko and Rado settled it for $n\geqslant n_{0}(k)$ in the famous Erdős–Ko–Rado theorem [erdos1961intersection]. While Problem 2 was completely solved by Ahlswede and Khachatrian [ahlswede1997complete] after several improvements due to Frankl [frankl1978erdos], Wilson [wilson1984exact] and Frankl and Füredi [frankl1991beyond], Problem 1 remains open. Erdős and Sós conjectured [^1] The original conjecture of Erdős and Sós is for families avoiding intersection size $1$, the general case is due to Erdős. that any extremal solution of Problem 1 has the following form:
@@
for some set $T$ of size $t$, provided $k\geqslant 2t$ and $n\geqslant n_{0}(k)$. [^2] Note that $\mathcal{F}_{T}$ is the unique up to an isomorphism extremal example for Problem 2 if $n\geqslant(t+1)(k-t+1)$ [frankl1978erdos, wilson1984exact]. The conjecture was verified by Frankl and Füredi in the seminal paper [frankl_forbidding_1985]. However, the other regimes of parameters, when $t$ and $k$ grow with $n$, are also important and found numerous applications in discrete geometry [frankl1990partition, kahn1993counterexample], communication complexity [sgall1999bounds] and quantum computing [buhrman1998quantum]. Great progress on the problem was made by Keller and Lifshitz who solved Problem 1 when $n$ is linear in $k$ [keller2021junta], i.e. $n\geqslant f(t)k$ for some function $f(\cdot)$. Later, Ellis, Keller and Lifshitz [ellis2024stability] extended the previous result to the regime when $k\leqslant(1/2-\varepsilon)n$ for any small enough $\varepsilon$ provided $n\geqslant n_{0}(t,\varepsilon)$. Recently, Kupavskii and Zakharov [kupavskii_spread_2022] managed to solve Problem 1 under the polynomial dependencies between $n,k,t$, e.g. when $n=k^{\alpha}$, $t=k^{\beta}$ with $0<\beta<1/2$ and $\alpha>1+2\beta$, provided $k\geqslant k_{0}(\alpha,\beta)$. Partial results were obtained by Cherkashin [cherkashin2024set] and Keevash, Mubayi and Wilson [keevash2006set]. It is also worth mentioning the celebrated Frankl and Wilson result [frankl1981intersection] that bounds the size of $\mathcal{F}$ in Problem 1 by $\binom{n}{k-t}$ provided $k\geqslant 2t-1$ and kt+1k-t+1 is a prime power.
In recent decades, a rich and active direction of research has emerged: extending results from extremal set theory to families of various mathematical objects. The aforementioned Erdős–Ko–Rado theorem [erdos1961intersection] has received numerous extensions to families of permutations [frankl1977maximum, cameron2003intersecting, larose2004stable, ellis2011intersecting, kupavskii_spread_2022, keller_t-intersecting_2024, kupavskii2024almost], finite vector spaces [frankl1986erdos], integer sequences [ahlswede1998diametric, frankl_erdos-ko-rado_1999], matchings [kamat2013erdHos, fuentes2025intersecting], independent sets of a graph [borg2009erdHos, borg2014erdos, hurlbert2011erdHos, woodroofe2011erdHos, frankl2024holroyd](see also the survey [hurlbert2025survey]) and many other objects [chvatal2006intersecting, frankl2025intersecting, saengrungkongka2025t, iarovikova2025complete, kupavskii2023erd, kupavskii2023intersection](see also the survey [ellis2022intersection]).
In contrast, there are much fewer works on generalization of Problem 1 to other domains, mostly due to its hardness even in the $\binom{[n]}{k}$ case. There are some results for permutations [ellis2014forbidding, ellis2022approximation, keevash2017frankl], vector spaces [ellis2023forbidden] and integer sequences [keevash2023forbidden]. Significant progress in this direction was made by Kupavskii and Zakharov [kupavskii_spread_2022]. In particular, they present results for subfamilies of $m]^{n},\binom{[n]}{k}^{w}$, permutations and $(n,k,w)$-designs. However, assumptions of their theory turn out to be quite restrictive in our setup, which we discuss next.
@@
Note that for $t=1$ this problem was resolved independently by Frankl and Tokushige [EKR_for_codes] and Ahlswede and Khachatrian [ahlswede1998diametric]. A more general result was obtained by Keevash and coauthors [keevash2023forbidden].
Theorem 1 (Theorem 1.1 of [keevash2023forbidden]).
@@
Another result was obtained by Kupavskii and Zakharov [kupavskii_spread_2022].
Theorem 2 (Proposition 23 of [kupavskii_spread_2022]).
@@
The proof combines ideas from Keevash and coauthors [keevash2023forbidden], Kupavskii and Zakharov [kupavskii_spread_2022], some recent developments of Kupavskii [kupavskii2023intersection], a paper of Keller and coauthors on hypercontractivity for some pseudorandom functions [keller_sharp_2023] and a paper of Kupavskii and the second author [kupavskii2025lineardependenciespolynomialfactors].
Our proof technique can be adapted to Problem 1, yielding the solution when $k\geqslant\operatorname{poly}(t)$ and $n\geqslant\operatorname{poly}(t)\cdot k$, but we postpone it to more general theory [kupavskii_noskov_followup].
@@ 3 Preliminaries
We treat the case $m\leqslant e^{Cn/t^{2}}$ using a measure boosting argument, which dates back to Dinur and Friedgut [dinur_intersecting_2009]. We illustrate it on subfamilies of sets from some ambient family $\mathcal{A}$ (which is $\mathcal{C}$ in our case) endowed with the uniform measure $\mu(\cdot)$. The argument is based on the following informal observation: given two pseudorandom families $\mathcal{F}_{1},\mathcal{F}_{2}$, under mild assumptions on their measure $\mu(\cdot)$ in some ambient family $\mathcal{A}$, one can obtain families $\widetilde{\mathcal{F}}_{1},\widetilde{\mathcal{F}}_{2}$ in (probably, different) ambient family $\widetilde{\mathcal{A}}$ such that
@@
The efficiency of this idea significantly varies with the notion of pseudorandomness and the range family $\widetilde{\mathcal{A}}$ used. The original definition of pseudorandomness due to Dinur and Friedgut called uncapturability plays the key role in the celebrated junta method [keller2021junta, ellis2023forbidden, keevash2023forbidden, lifshitz2020set]. Later, Alweiss and coathours [alweiss2021improved] made a breakthrough result in the Erdős sunflower problem using another notion of pseudorandomness called spreadness. Using their results together with delicate combinatorial arguments, Kupavskii and Zakharov [kupavskii_spread_2022] developed the spread approximation technique for which they introduced a new concept of pseudorandomness named $\tau$-homogeneity. They apply their method to obtain solutions of the Deza–Frankl problem on $t$-intersecting permutations and the Erdős–Sós forbidden intersection problem under polynomial dependencies between parameters. The technique was further improved in subsequent works of Kupavskii and coathours [kupavskii2023erd, kupavskii2023intersection, kupavskii2024almost, kupavskii2025lineardependenciespolynomialfactors, frankl2025hajnal, iarovikova2025complete] yielding a large variety of results in extremal set theory. In particular, in this paper we enhance the spread approximation technique for the families of codes. At the same time, Keevash and coathours [keevash2021global, keevash2023forbidden] presented the notion of global families, and using it they reached large progress in Turán-type problems on forbidden large sparse hypergraphs. Then, independently of Kupavskii and Zakharov, they modified their definition of global families (so it coincides now with the $\tau$-homogeneity), and generalized it for functions on a product space [keller_sharp_2023]. The main result of their findings is a new hypercontractivity inequality for global functions, which we use below; see Theorem 10. In follow-up papers, they show that their approach has a lot of consequences in different fields of discrete mathematics [gill2024initiating, lifshitz2024completing, keller2024improved, filmus2024hypercontractivity].
@@ 3.1 Random gluings
Definition 1 (Definition 6.5 [keevash2023forbidden]).
@@
Definition 2 (Definition 6.7 [keevash2023forbidden]).
@@ 3.1.1 Gluings and the measure boosting argument
Claim 5 (Claim 6.8 [keevash2023forbidden]).
@@
$(1-\lambda_{*}(\widetilde{T}))^{2}=\sup\{\mathbb{E}_{\mathbf{x}\sim\nu}(\widetilde{T}f(\mathbf{x}))^{2}\;:\;\mathbb{E}_{\mathbf{x}\sim\nu}f(\mathbf{x})=0,\;\mathbb{E}_{\mathbf{x}\sim f}f^{2}(\mathbf{x})=1\}.$
@@
For the proof, we refer the reader to that of Lemma 6.9 [keevash2023forbidden].
@@
Lemma 7 (Lemma 5.13[keevash2023forbidden]).
@@
Proposition 9 (Lemma 5.11 [keevash2023forbidden]).
@@
To control $\operatorname{Stab}_{\rho_{0}}(\mathcal{F})$, one need to guarantee some pseudorandomness of $\mathcal{F}$. Authors of the original paper [keevash2023forbidden] used the notion of globalness due to Keevash et al. [keevash2021global]. Instead, we use the concept of globalness from [keller_sharp_2023].
Theorem 10 (Theorem 1.4 from [keller_sharp_2023]).
@@
In the paper [keller_sharp_2023], the proof of Theorem 10 is presented for the case when all $\Omega_{i}$ and $\nu_{i}$ coincide. However, it works for the product space since it based on the freezing argument and induction.
@@
Lemma 11 (Lemma 5.9 [keevash2023forbidden]).
@@
There is an alternative measure boosting argument that performs well if one of the initial measures $\mu(\mathcal{F}_{1}),\mu(\mathcal{F}_{2})$ is large. We present it as the following lemma. While its statement can not be found in paper [keevash2023forbidden], the proof coincides with that of Theorem 2.3 [keevash2023forbidden]. For the sake of completeness, we provide the proof in the appendix.
@@ 3.2 Homogeneous families
Lemma 14 (Lemma 52 from [kupavskii2025lineardependenciespolynomialfactors]).
@@
In the paper [kupavskii2025lineardependenciespolynomialfactors], the authors require $\mu(\mathcal{R})\leqslant 32\tau^{-q}$, but their proof works with the tighter bound suggested in the above lemma. For the sake of completeness, let us present the proof of Lemma 14.
@@
Assume for a while, that pairs $(Z_{i},x_{i})$ obtained by the spread approximation lemma 14 are $t$-agreeing, i.e. for any $(Z_{i},x_{i}),(Z_{j},x_{j})$ from $\{(Z_{i},x_{i})\}_{i=1}^{l}$, we have $\operatorname{agr}_{{Z_{i}\cap Z_{j}}}(x_{i},x_{j})\geqslant t$. In the next section, we discuss a number of arguments due to papers [kupavskii_spread_2022] and [kupavskii2025lineardependenciespolynomialfactors], which help to deal with this and weaker properties of the spread approximations and to prove that there exists a pair $(T,x)$, $|T|=t$ and $x\in[m]^{T}$, such that for almost all codes $y\in\mathcal{F}$, we have $y|_{T}=x$.
@@ 3.3 Simplification arguments
Simplification arguments are one of the key components of the spread approximation technique, but they vary from problem to problem [kupavskii_spread_2022, kupavskii2023erd, kupavskii2024almost, frankl2025hajnal, kupavskii2025lineardependenciespolynomialfactors, iarovikova2025complete]. All of them are designed for families of sets from an ambient family $\mathcal{A}$ that meet some conditions. First, we list them.
@@
Then, we require the following assumption, which is a stronger version of Assumption 1 [kupavskii2025lineardependenciespolynomialfactors].
@@
are $t$-agreeing, then the corresponding sets $\widetilde{Z}_{i}\in\partial_{\leqslant q}\mathcal{C}$, $\widetilde{Z}_{i}\cap I_{j}=\{(j-1)m+(x^{j})_{i}\}$ for $j\in Z_{i}$ and $i=1,\ldots,l$, are $t$-intersecting. We say that a $t$-intersecting family $\widetilde{\mathcal{S}}=\{\widetilde{Z}_{i}\}$ is trivial, if there exists a set $\widetilde{T}$ of size $t$ such that $\widetilde{T}\subset\widetilde{Z}$ for all $\widetilde{Z}\in\widetilde{\mathcal{S}}$. It turns out that any non-trivial $t$-intersecting family of restrictions spans a much smaller number of sets than a single set of size $t$. This was shown in the paper [kupavskii_spread_2022] under Assumption 1. We use the following improved result due to [kupavskii2023erd].
Theorem 19 (Theorem 14 [kupavskii2023erd]).
@@
Next, we move on to the discussion of the regime $m\geqslant n^{Ct}$ and tools required for its analysis, which were developed in the paper of Kupavskii and the second author [kupavskii2025lineardependenciespolynomialfactors]. First, we give some prelimanries concerning sunflowers and $(\mathcal{S},s,t)$-systems. Following Definition 8 of [kupavskii2025lineardependenciespolynomialfactors], we will name a family $\mathcal{S}=\{(Z_{i},x_{i})\}_{i=1}^{l}$ obtained from the spread approximation in Lemma 14.
@@
To tackle the case $m\geqslant n^{Ct}$, we will adapt the technique of $(\mathcal{S},s,t)$-systems of [kupavskii2025lineardependenciespolynomialfactors] which was developed for systems without sunflowers. Hence, we need to introduce the definition of a sunflower.
@@
It is easy to check that our problem corresponds to forbidding a sunflower with $2$ petals and the core of size t1t-1 in a subfamily of $\mathcal{C}$. Hence, we may apply the main results of [kupavskii2025lineardependenciespolynomialfactors] on $(\mathcal{S},s,t)$-systems:
@@
Lemma 21 (Lemma 56 of [kupavskii2025lineardependenciespolynomialfactors]).
@@
Lemma 22 (Proposition 57 of [kupavskii2025lineardependenciespolynomialfactors]).

### v2

@@ Abstract
In this paper, we study the famous Erdős–Sós forbidden intersection problem for words over an alphabet of size $m$: what is the maximal size of a subfamily $\mathcal{F}$ of $m]^{n}$ that does not contain two vectors $x,y$ coinciding on exactly t1t-1 coordinates? We answer this question provided $m\geqslant\operatorname{poly}(t)$ and $n\geqslant\operatorname{poly}(t)$ for some polynomial function $\operatorname{poly}(\cdot)$ of $t$, greatly extending the recent result of Keevash, Lifshitz, Long and Minzer. Our proof combines some of the recently developed methods in extremal combinatorics, including the spread approximation technique of Kupavskii and Zakharov and the hypercontractivity approach developed in a series of works by Keevash, Keller, Lifshitz, Long, Marcus and Minzer.
@@ 1 Introduction
Let $\binom{[n]}{k}$ be a family of $k$-subsets of $n]=\{1,\ldots,n\}$. In 1971, Erdős and Sós (see [19]) considered the following problem.
@@
Problem 1 is a harder version of the following question raised by Erdős, Ko and Rado [18] ten years earlier.
@@
Erdős, Ko and Rado settled it for $n\geqslant n_{0}(k)$ in the famous Erdős–Ko–Rado theorem [18]. While Problem 2 was completely solved by Ahlswede and Khachatrian [1] after several improvements due to Frankl [33], Wilson [63] and Frankl and Füredi [24], Problem 1 remains open. Erdős and Sós conjectured [^1] The original conjecture of Erdős and Sós is for families avoiding intersection size $1$, the general case is due to Erdős. that any extremal solution of Problem 1 has the following form:
@@
for some set $T$ of size $t$, provided $k\geqslant 2t$ and $n\geqslant n_{0}(k)$. [^2] Note that $\mathcal{F}_{T}$ is the unique up to an isomorphism extremal example for Problem 2 if $n\geqslant(t+1)(k-t+1)$ [33, 63]. The conjecture was verified by Frankl and Füredi in the seminal paper [23]. However, the other regimes of parameters, when $t$ and $k$ grow with $n$, are also important and found numerous applications in discrete geometry [28, 39], communication complexity [62] and quantum computing [6]. Great progress on the problem was made by Keller and Lifshitz who solved Problem 1 when $n$ is linear in $k$ [49], i.e. $n\geqslant f(t)k$ for some function $f(\cdot)$. Later, Ellis, Keller and Lifshitz [14] extended the previous result to the regime when $k\leqslant(1/2-\varepsilon)n$ for any small enough $\varepsilon$ provided $n\geqslant n_{0}(t,\varepsilon)$. Recently, Kupavskii and Zakharov [52] managed to solve Problem 1 under the polynomial dependencies between $n,k,t$, e.g. when $n=k^{\alpha}$, $t=k^{\beta}$ with $0<\beta<1/2$ and $\alpha>1+2\beta$, provided $k\geqslant k_{0}(\alpha,\beta)$. Partial results were obtained by Cherkashin [8] and Keevash, Mubayi and Wilson [44]. It is also worth mentioning the celebrated Frankl and Wilson result [31] that bounds the size of $\mathcal{F}$ in Problem 1 by $\binom{n}{k-t}$ provided $k\geqslant 2t-1$ and kt+1k-t+1 is a prime power.
In recent decades, a rich and active direction of research has emerged: extending results from extremal set theory to families of various mathematical objects. The aforementioned Erdős–Ko–Rado theorem [18] has received numerous extensions to families of permutations [21, 7, 56, 12, 52, 46, 55], finite vector spaces [32], integer sequences [2, 29], matchings [40, 34], independent sets of a graph [4, 5, 36, 64, 26](see also the survey [37]) and many other objects [9, 25, 61, 38, 53, 54](see also the survey [17]).
In contrast, there are much fewer works on generalization of Problem 1 to other domains, mostly due to its hardness even in the $\binom{[n]}{k}$ case. There are some results for permutations [16, 15, 43, 48], vector spaces [11] and integer sequences [42]. Significant progress in this direction was made by Kupavskii and Zakharov [52]. In particular, they present results for subfamilies of $m]^{n},\binom{[n]}{k}^{w}$, permutations and $(n,k,w)$-designs. However, assumptions of their theory turn out to be quite restrictive in our setup, which we discuss next.
@@
Note that for $t=1$ this problem was resolved independently by Frankl and Tokushige [30] and Ahlswede and Khachatrian [2]. A more general result was obtained by Keevash and coauthors [42].
Theorem 1 (Theorem 1.1 of [42]).
@@
Another result was obtained by Kupavskii and Zakharov [52].
Theorem 2 (Proposition 23 of [52]).
@@
The proof combines ideas from Keevash and coauthors [42], Kupavskii and Zakharov [52], some recent developments of Kupavskii [54], a paper of Keller and coauthors on hypercontractivity for some pseudorandom functions [45] and a paper of Kupavskii and the second author [50].
Our proof technique can be adapted to Problem 1, yielding the solution when $k\geqslant\operatorname{poly}(t)$ and $n\geqslant\operatorname{poly}(t)\cdot k$, but we postpone it to more general theory [51].
@@ 3 Preliminaries
We treat the case $m\leqslant e^{Cn/t^{2}}$ using a measure boosting argument, which dates back to Dinur and Friedgut [10]. We illustrate it on subfamilies of sets from some ambient family $\mathcal{A}$ (which is $\mathcal{C}$ in our case) endowed with the uniform measure $\mu(\cdot)$. The argument is based on the following informal observation: given two pseudorandom families $\mathcal{F}_{1},\mathcal{F}_{2}$, under mild assumptions on their measure $\mu(\cdot)$ in some ambient family $\mathcal{A}$, one can obtain families $\widetilde{\mathcal{F}}_{1},\widetilde{\mathcal{F}}_{2}$ in (probably, different) ambient family $\widetilde{\mathcal{A}}$ such that
@@
The efficiency of this idea significantly varies with the notion of pseudorandomness and the range family $\widetilde{\mathcal{A}}$ used. The original definition of pseudorandomness due to Dinur and Friedgut called uncapturability plays the key role in the celebrated junta method [49, 11, 42, 57]. Later, Alweiss and coathours [3] made a breakthrough result in the Erdős sunflower problem using another notion of pseudorandomness called spreadness. Using their results together with delicate combinatorial arguments, Kupavskii and Zakharov [52] developed the spread approximation technique for which they introduced a new concept of pseudorandomness named $\tau$-homogeneity. They apply their method to obtain solutions of the Deza–Frankl problem on $t$-intersecting permutations and the Erdős–Sós forbidden intersection problem under polynomial dependencies between parameters. The technique was further improved in subsequent works of Kupavskii and coathours [53, 54, 55, 50, 27, 38] yielding a large variety of results in extremal set theory. In particular, in this paper we enhance the spread approximation technique for the families of codes. At the same time, Keevash and coathours [41, 42] presented the notion of global families, and using it they reached large progress in Turán-type problems on forbidden large sparse hypergraphs. Then, independently of Kupavskii and Zakharov, they modified their definition of global families (so it coincides now with the $\tau$-homogeneity), and generalized it for functions on a product space [45]. The main result of their findings is a new hypercontractivity inequality for global functions, which we use below; see Theorem 10. In follow-up papers, they show that their approach has a lot of consequences in different fields of discrete mathematics [35, 58, 47, 20].
@@ 3.1 Random gluings
Definition 1 (Definition 6.5 [42]).
@@
Definition 2 (Definition 6.7 [42]).
@@ 3.1.1 Gluings and the measure boosting argument
Claim 5 (Claim 6.8 [42]).
@@
$(1-\lambda_{*}(T))^{2}=\sup\{\mathbb{E}_{\mathbf{x}\sim\nu}(Tf(\mathbf{x}))^{2}\;:\;\mathbb{E}_{\mathbf{x}\sim\nu}f(\mathbf{x})=0,\;\mathbb{E}_{\mathbf{x}\sim f}f^{2}(\mathbf{x})=1\}.$
@@
For the proof, we refer the reader to that of Lemma 6.9 [42].
@@
Lemma 7 (Lemma 5.13[42]).
@@
Proposition 9 (Lemma 5.11 [42]).
@@
To control $\operatorname{Stab}_{\rho_{0}}(\mathcal{F})$, one need to guarantee some pseudorandomness of $\mathcal{F}$. Authors of the original paper [42] used the notion of globalness due to Keevash et al. [41]. Instead, we use the concept of globalness from [45].
Theorem 10 (Theorem 1.4 from [45]).
@@
In the paper [45], the proof of Theorem 10 is presented for the case when all $\Omega_{i}$ and $\nu_{i}$ coincide. However, it works for the product space since it based on the freezing argument and induction.
@@
Lemma 11 (Lemma 5.9 [42]).
@@
There is an alternative measure boosting argument that performs well if one of the initial measures $\mu(\mathcal{F}_{1}),\mu(\mathcal{F}_{2})$ is large. We present it as the following lemma. While its statement can not be found in paper [42], the proof coincides with that of Theorem 2.3 [42]. For the sake of completeness, we provide the proof in the appendix.
@@ 3.2 Homogeneous families
Lemma 14 (Lemma 52 from [50]).
@@
In the paper [50], the authors require $\mu(\mathcal{R})\leqslant 32\tau^{-q}$, but their proof works with the tighter bound suggested in the above lemma. For the sake of completeness, let us present the proof of Lemma 14.
@@
Assume for a while, that pairs $(Z_{i},x_{i})$ obtained by the spread approximation lemma 14 are $t$-agreeing, i.e. for any $(Z_{i},x_{i}),(Z_{j},x_{j})$ from $\{(Z_{i},x_{i})\}_{i=1}^{l}$, we have $\operatorname{agr}_{{Z_{i}\cap Z_{j}}}(x_{i},x_{j})\geqslant t$. In the next section, we discuss a number of arguments due to papers [52] and [50], which help to deal with this and weaker properties of the spread approximations and to prove that there exists a pair $(T,x)$, $|T|=t$ and $x\in[m]^{T}$, such that for almost all codes $y\in\mathcal{F}$, we have $y|_{T}=x$.
@@ 3.3 Simplification arguments
Simplification arguments are one of the key components of the spread approximation technique, but they vary from problem to problem [52, 53, 55, 27, 50, 38]. All of them are designed for families of sets from an ambient family $\mathcal{A}$ that meet some conditions. First, we list them.
@@
Then, we require the following assumption, which is a stronger version of Assumption 1 [50].
@@
are $t$-agreeing, then the corresponding sets $\widetilde{Z}_{i}\in\partial_{\leqslant q}\mathcal{C}$, $\widetilde{Z}_{i}\cap I_{j}=\{(j-1)m+(x^{j})_{i}\}$ for $j\in Z_{i}$ and $i=1,\ldots,l$, are $t$-intersecting. We say that a $t$-intersecting family $\widetilde{\mathcal{S}}=\{\widetilde{Z}_{i}\}$ is trivial, if there exists a set $\widetilde{T}$ of size $t$ such that $\widetilde{T}\subset\widetilde{Z}$ for all $\widetilde{Z}\in\widetilde{\mathcal{S}}$. It turns out that any non-trivial $t$-intersecting family of restrictions spans a much smaller number of sets than a single set of size $t$. This was shown in the paper [52] under Assumption 1. We use the following improved result due to [53].
Theorem 19 (Theorem 14 [53]).
@@
Next, we move on to the discussion of the regime $m\geqslant n^{Ct}$ and tools required for its analysis, which were developed in the paper of Kupavskii and the second author [50]. First, we give some prelimanries concerning sunflowers and $(\mathcal{S},s,t)$-systems. Following Definition 8 of [50], we will name a family $\mathcal{S}=\{(Z_{i},x_{i})\}_{i=1}^{l}$ obtained from the spread approximation in Lemma 14.
@@
To tackle the case $m\geqslant n^{Ct}$, we will adapt the technique of $(\mathcal{S},s,t)$-systems of [50] which was developed for systems without sunflowers. Hence, we need to introduce the definition of a sunflower.
@@
It is easy to check that our problem corresponds to forbidding a sunflower with $2$ petals and the core of size t1t-1 in a subfamily of $\mathcal{C}$. Hence, we may apply the main results of [50] on $(\mathcal{S},s,t)$-systems:
@@
Lemma 21 (Lemma 56 of [50]).
@@
Lemma 22 (Proposition 57 of [50]).